Julienne deposited her money amounting to P5,500 to her saving account in the CCJE cooperative. If the coop gives 6% interest per annum. How much would her money be after 10 years?
Found 2 solutions by ikleyn, ankor@dixie-net.com:
Answer by ikleyn(52784)

.

To solve the problem,  we do not need to know the name of this banking institution.

What we  REALLY  need to know is if the account is compounded or simple.

It is  90%  probability that it is compounded,  but the problem,  had it be formulated  PROFESSIONALLY,
should say it  directly,  openly  and  explicitly,  without leaving the room for the reader for his  (or her)  guessing or assumptions.



Answer by ankor@dixie-net.com(22740)

Julienne deposited her money amounting to P5,500 to her saving account in the CCJE cooperative.
If the coop gives 6% interest per annum.
How much would her money be after 10 years?
:
Assume it is compounded annually A = Ao(1+r)^t, where
A = amt after t time
Ao = initial amt
r = interest rate in decimal form
t = time in yrs
:
A = 500(1 + .06)^10
A = 500(1.06)^10
Find 1.06^10 on your calc
A = 500 * 1.7980477
A = $895.42 after 10 yrs